X Company is considering modifying one of its main products next year that would make it more attractive to customers. Financial information for the product for this year was as follows:

Selling price $9.32
Variable costs per unit $3.54
Total fixed costs $10,220


If it modified the product, X Company could increase the selling price by $2.10, and unit sales for the modified product would still be the same as current product unit sales. However, fixed costs would increase to $13,770. At what unit sales level would X Company be indifferent between modifying the product and not modifying it?
Answer #1

Calculate indifferent point

(9.32-3.54)X-10220 = (11.42-3.54)X-13770

5.78X-10220 = 7.88-13770

-2.10X = -3550

X(Indifferent point) = 3550/2.1 = 1690 Units